OpenStudy (anonymous):

What is the equivalent of pi over 3 radians in degrees? 30 degrees 45 degrees 60 degrees 90 degrees

OpenStudy (anonymous):

@mathway I have this and 1 more :)

OpenStudy (anonymous):

π/3×180°/π = 60°

OpenStudy (anonymous):

^ do what s/he said

OpenStudy (anonymous):

Conversion from Radian to Degrees; \[\frac{ r }{ \pi } = \frac{ d }{ 180 }\] In your case, r would be \[r = \frac{ \pi }{ 3 }\] Putting it into the formula; \[\frac{ \frac{ \pi }{ 3 } }{ \pi } = \frac{ d }{ 180 }\] Turn it around to get d on one side (degree); \[d = \frac{ \frac{ \pi }{ 3 } * 180 }{ \pi } \] Which means; \[d = 60\] Note that you can use the same formula in order to get the radian - just input your degree and solve for r. Hope it helps.

OpenStudy (anonymous):

Thank you! I have one more that i've been trying to get help with for like ever now, Can you help on that?

OpenStudy (anonymous):

yeah sure, just open a new question!